If the line l1 and the line y = -x+2 are symmetric about the y-axis, the equation of the line l1 is...?
Found 2 solutions by math_helper, greenestamps:
Answer by math_helper(2461)


If y=f(x) is a line, and you wish to obtain the reflection about the y-axis, you replace "x" with "-x" everywhere:
y = f(x) = -x+2
so the reflection about the y-axis is:
f(-x) = -(-x)+2 = x+2

Answer by greenestamps(13200)



The response from the other tutor shows a formal way of finding the answer.

On a less formal and more intuitive level, if the two lines are symmetric about the y-axis then their slopes are the same size but opposite signs, and the y-intercepts are the same.

The given line y = -x+2 has slope -1 and y-intercept 2, so the second line has slope 1 and y-intercept 2

ANSWER: y = x+2